A particle, with a +2uC charge and a kinetic energy of 0.1 Joules, is dispersed in a uniform magnetic field of 0.1 Tesla. If the particle moves in a circular trajectory with a radius of 2m, what is the particle's velocity? Answer:5×10^5 m/s
Answer #1

kE = 0.5 m v^2

0.5* ( m v) v = 0.1

(mv) = 0.2 / v

now as we know

r = mv / qB

2 = (0.2 / v) / (2* 10^-6* 0.1)

v = 5* 10^5 m/s

=====

Comment before rate in case any doubt, will reply for sure.. goodluck